OpenStudy (tatum0725):

Rhombus with a perimeter of 20 meter and diagonal of 8 meters

OpenStudy (minecraftgeek5):

umm is this a question it seems very incomplete

OpenStudy (tatum0725):

Finding the area . Rhombus with a perimeter of 20 meter and diagonal of 8 meter

OpenStudy (matlee):

if you search this on google : area of a rhombus calculator It will tell you the formula so if you want to work it out i can help

OpenStudy (mahoganie.carson):

All four sides are equal, so each side is 5 m. Also the diagonals right-bisect each other to form 4 congruent right-angled triangles Consider one of thes. The hypotenuse is 5, one leg is 4 , so the other must be 3 by Pythagoras. Area of one of those = (1/2)(3)(4) = 6 So the whole rhombus has area of 24 m^2 or the area of a rhomus = (1/2)(product of their diagonals) = (1/2)(6)(8) = 24 :)